武谷三男博士の業績リスト（第3版）

矢野 忠^{1}

1 はじめに

すでにこの「素粒子論研究」に第 1 版 [1], 第 2 版 [2] を発表したが, 武谷の年譜を作成しようとすれば，論文の発行の年月までリストに記述をしておいた方がよいことがわかった．また１つだけ新しい論文を見つけた．それでここに３度目の改訂をする．この第 3 版の収録データは (1) 著者名, (2) 論文名, (3) 雑誌名または書名,

(4) 巻号または号, 発行年月, 頁, である. 収録データについて注記をしておきたい. 掲載順は発行順を原則としたが, 学術雑誌は発行年または発行年月しか分からないものはその発行年または発行年月の末尾に加えた. また, たとえば Prog. Thoer. Phys. (P. T. P. と略す）に掲載された論文がその P. T. P. Supplement (Suppl. と略す) に再掲されているのも重複して載せている. また Suppl. No. 50 に日本語からの英訳として出ているもので, 狭い意味では物理の研究論文には入らないかもしれないものや Suppl. No. 105 のインタビュー等も載せている. なお, 武谷の著作については [3], [4], [5], [6] および [7] を参照されたい.

(2010.6.22)

2 List of the Scientific Papers of Mituo Taketani

1. M. Taketani A Collection of Land Mollusca from Taiwan (Formosa) Transactions of the National History Society of Formosa, Vol. 26 No. 154 (1936) 271-278.

2. H. Yukawa, S. Sakata and M. Taketani On the Interaction of Elementary Particles. III Proc. Phys.-Math. Soc. Japan, 3rd Series Vol. 20 No. 3 (Mar. 1938) 319-340.

3. H. Yukawa, S. Sakata, M. Kobayasi and M. Taketani On the Interaction of Elementary Particles. IV Proc. Phys.-Math. Soc. Japan, 3rd Series Vol. 20 No. 8 (Aug. 1938) 720-745.

4. S. Sakata and M. Taketani Versuch einer Theorie des β-Zerfalls Proc. Phys.-Math. Soc. Japan, 3rd Series Vol. 20 No. 10 (Oct. 1938) 962-963.

5. M. Taketani and S. Sakata On the Wave Equation of Meson Proc. Phys.-Math. Soc. Japan, 3rd Series Vol. 22 No. 8 (Aug. 1940) 757-770.

6. S. Sakata and M. Taketani Note on Casimir’s Method of the Spin Summation in the Case of the Meson Sci. Pap. I. P. C. R., Vol. 38 No. 984 (Sep. 1940) 1-11.

1e-mail: yanotad@earth.ocn.ne.jp

7. M. Taketani and S. Sakata On the Wave Equation of Meson, Errata Proc. Phys.-Math. Soc. Japan, 3rd Series Vol. 23 No. 1 (Jan. 1941) 160.

8. M. Taketani Conflict between Matter and Field

- An Analysis of the Difficulties of the Theory of Elementary Particles - Prog. Theor. Phys., Vol. 2 No. 4 (Nov.-Dec. 1947) 187-197.

9. M. Taketani On the Neutral Meson Prog. Theor. Phys., Vol. 3 No. 4 (Oct.-Dec. 1948) 349-355.

10. M. Taketani, S. Nakamura, K. Ono and M. Sasaki On the Two-Meson Theory Phys. Rev., 2nd Series Vol. 76 No. 1 (Jul. 1949) 60-63.

11. M. Taketani, S. Nakamura and M. Sasaki On the μ-Meson Capture Phys. Rev., 2nd Series Vol. 76 No. 6 (Sep. 1949) 852.

12. M. Taketani, M. Sasaki and S. Nakamura On the mu-Meson Capture Prog. Theor. Phys., Vol. 4 No. 4 (Oct.-Dec. 1949) 552-556.

13. M. Taketani, S. Machida and T. Tamura On the Electromagnetic Properties of Nucleons Prog. Theor. Phys., Vol. 5 No. 1 (Jan.- Feb. 1950) 150-152.

14. M. Taketani, S. Nakamura and M. Sasaki On the Method of the Theory of Elementary Particles Prog. Theor. Phys., Vol. 5 No. 4 (Jul.-Aug. 1950) 730-734.

15. S. Nakamura, H. Fukuda, K. Ono, M. Sasaki and M. Taketani Analysis on the Two Meson Theory Prog. Theor. Phys., Vol. 5 No. 4 (Jul.-Aug. 1950) 740-747.

16. M. Taketani, S. Nakamura, M. Umezawa, K. Ono and Y. Yamaguchi ft-Values of β-Decay and the Shell Models Phys. Rev., Vol. 82 No. 3 (Mar. 1951) 441-442.

17. M. Taketani, S. Nakamura, K. Ono and M. Umezawa On the Matrix Elements of the Beta Decay Prog. Theor. Phys., Vol. 6 No. 3 (May-Jun. 1951) 286-294.

18. M. Umezawa, S. Nakamura, Y. Yamaguchi and M. Taketani Nuclear Shell Model and the β-Decay Schems, I Prog. Theor. Phys., Vol. 6 No. 3 (May-Jun. 1951) 408-425.

19. M. Taketani and S. Machida On the Production of Negative Protons Prog. Theor. Phys., Vol. 6 No. 4 (Jul.-Aug. 1951) 559-571.

20. M. Taketani, S. Nakamura and M. Sasaki On the Method of the Theory of Nuclear Forces Prog. Theor. Phys., Vol. 6 No. 4 (Jul.-Aug. 1951) 581-586.

21. M. Taketani, S. Onuma and S. Koide On the Nuclear Forces and the Ground State of Deuteron Prog. Theor. Phys., Vol. 6 No. 4 (Jul.-Aug. 1951) 635-637.

22. M. Taketani, S. Machida and S. Onuma Meson Theory of Nuclear Forces Prog. Theor. Phys., Vol. 6 No. 4 (Jul.-Aug. 1951) 638-639.

23. S. Machida, S. Onuma and M. Taketani On the Fourth-Order Nuclear Potential Prog. Theor. Phys., Vol. 6 No. 5 (Sep.-Oct. 1951) 904-905.

24. M. Taketani, S. Machida and S. Onuma The Meson Theory of Nuclear Forces, I

- The Deuteron Ground State and Low Energy Neutron-Proton Scattering - Prog. Theor. Phys., Vol. 7 No. 1 (Jan. 1952) 45-56.

25. S. Fujii, J. Iwadare, S. Otsuki, M. Taketani, S. Tani and W. Watari Note on the Meson Theory of Nuclear Force Prog. Theor. Phys., Vol. 10 No. 4 (Oct. 1953) 478-480.

26. M. Taketani Studies on Nuclear Forces in Japan Proceedings of the International Conference of Theoretical Physics, (Sep. 1953), Kyoto and Tokyo, 241-244.

27. S. Fujii, J. Iwadare, S. Otsuki, M. Taketani, S. Tani and W. Watari Meson Theory of Nuclear Forces, II - High Energy Nucleon-Nucleon Scattering - Prog. Theor. Phys., Vol. 11 No. 1 (Jan. 1954) 11-19.

28. N. Fukuda, K. Sawada and M. Taketani On the Construction of Potential in Field Theory Prog. Theor. Phys., Vol. 12 No. 2 (Aug. 1954) 156-166.

29. H. Takebe, S. Nakamura and M. Taketani β-Ray Spectrum of RaE Prog. Theor. Phys., Vol. 14 No. 4 (Oct. 1955) 317-336.

30. H. Yukawa, S. Sakata and M. Taketani On the Interaction of Elementary Particles. III (Reprinted from Proc. Phys.-Math. Soc. Japan 20 (1938), 319-340) Collected Papers on Meson Theory I (Formalism and Models) Prog. Theor. Phys. Suppl., No. 1 (1955) 24-45.

31. H. Yukawa, S. Sakata, M. Kobayasi and M. Taketani On the Interaction of Elementary Particles. IV (Reprinted from Proc. Phys.-Math. Soc. Japan 20 (1938), 720-745) Collected Papers on Meson Theory I (Formalism and Models) Prog. Theor. Phys. Suppl., No. 1 (1955) 46-71.

32. M. Taketani and S. Sakata On the Wave Equation of Meson (Reprinted from Proc. Phys.-Math. Soc. Japan 22 (1940), 757-770) Collected Papers on Meson Theory I (Formalism and Models) Prog. Theor. Phys. Suppl., No. 1 (1955) 84-97.

33. S. Sakata and M. Taketani Note on Casimir’s Method of the Spin Summation in the Case of the Meson (Reprinted from Sci. Pap. I. P. C. R. 38 (1940), 1-11) Collected Papers on Meson Theory I (Formalism and Models) Prog. Theor. Phys. Suppl., No. 1 (1955) 98-108.

34. M. Taketani On the Neutral Meson (Reprinted from Prog. Theor. Phys. 3 (1948), 349-355) Collected Papers on Meson Theory I (Formalism and Models) Prog. Theor. Phys. Suppl., No. 1 (1955) 219-225.

35. M. Taketani, M. Sasaki and S. Nakamura On the mu-Meson Capture (Reprinted from Prog. Theor. Phys. 4 (1949), 552-556) Collected Papers on Meson Theory I (Formalism and Models) Prog. Theor. Phys. Suppl., No. 1 (1955) 226-230.

36. M. Taketani, S. Nakamura, K. Ono and M. Sasaki On the Two-Meson Theory (Reprinted from Phys. Rev. 76 (1949), 60-63.) Collected Papers on Meson Theory I (Formalism and Models) Prog. Theor. Phys. Suppl., No. 1 (1955) 234-237.

37. S. Nakamura, H. Fukuda, K. Ono, S. Sasaki and M. Taketani Analysis on the Two Meson Theory (Reprinted from Prog. Theor. Phys. 5 (1950), 740-747) Collected Papers on Meson Theory I (Formalism and Models) Prog. Theor. Phys. Suppl., No. 1 (1955) 244-251.

38. M. Taketani, T. Hatanaka and S. Obi Populations and Evolution of Stars Prog. Theor. Phys. Vol. 15 No. 2 (Feb. 1956) 89-94.

39. K. Inoue, S. Machida, M. Taketani and T. Toyoda Pion Theory of Nuclear Forces Prog. Theor. Phys., Vol. 15 No. 2 (Feb. 1956) 122-136.

40. M. Taketani Introduction. Nuclear Forces and Meson Theory Meson Theory III. Nuclear Forces Prog. Theor. Phys. Suppl., No. 3 (1956) 1-12.

41. M. Taketani, S. Nakamura and M. Sasaki Part V. On the Method of the Theory of Nuclear Forces (reprinted from Prog. Theor. Phys. 6 (1951), 581-586) Meson Theory III. Nuclear Forces Prog. Theor. Phys. Suppl., No. 3 (1956) 169-174.

42. S. Nakamura, T. Ohmura and M. Taketani A Note on the Nuclear Reactions in Stars Prog. Theor. Phys., Vol. 21 No. 1 (Jan. 1959) 18-28.

43. M. Taketani, Y. Katayama, P. L. Ferreira, G. W. Bund and P. R. de Paula e Silva Meson Decays and Structure of Muon Prog. Theor. Phys., Vol. 21 No. 6 (Jun. 1959) 799-817.

44. Y. Katayama, M. Taketani and J. L. Ferreira On the Origin of the Electron Mass Prog. Theor. Phys., Vol. 21 No. 6 (Jun. 1959) 818-826.

45. Y. Katayama, M. Taketani, S. Ragusa and D. R. de Oliveira On the Electromagnetic Structure of Nucleons and Their Mass Difference Prog. Theor. Phys., Vol. 23 No. 2 (Feb. 1960) 328-352.

46. M. Taketani and Y. Katayama On a Unified Description of Elementary Particles Prog. Theor. Phys., Vol. 24 No. 3 (Sep. 1960) 661-678.

47. M. Taketani and S. Machida On the Method of the Theory of Nuclear Forces Prog. Theor. Phys., Vol. 24 No. 6 (Dec. 1960) 1317-1324.

48. S. Otuki, M. Taketani, R. Tamagaki and W. Watari Pion Theory of Nuclear Forces and Low Energy P -Wave Phase Shifts Prog. Theor. Phys., Vol. 25 No. 3 (Mar. 1961) 427-435.

49. M. Taketani Chapter 2. A New Stage of Elementary Particle Theory Part I. Introduciton Structure of Elementary Particles Prog. Theor. Phys. Suppl., No. 19 (1961) 10-22 (pub. date: Mar. 1962).

50. Y. Katayama and M. Taketani Chapetr 4. On the Structure of Lepton Part III. Further Outlook for the Theory of Elementary Particles Structure of Elementary Particles Prog. Theor. Phys. Suppl., No. 19 (1961) 178-188 (pub. date: Mar. 1962).

51. M. Taketani and Y. Katayama On the Neutrino Unified Model for Elementary Particles Part III. Further Outlook for the Theory of Elementary Particles Structure of Elementary Particles Prog. Theor. Phys. Suppl., No. 19 (1961) 189-202 (pub. date: Mar. 1962).

52. Y. Hama, J. Osada and M. Taketani Ghost Scatterings Nuovo Cim., Vol. 25 No. 5 (Sep. 1962) 1054-1060.

53. M. Taketani and M. Sawamura On the Existence of a Boson μ^{0} Prog. Theor. Phys., Vol. 27 No. 4 (Oct. 1962) 1287-1289.

54. M. Taketani and T. Tati ΔS/ΔQ = −1 Decay and “ Induced Transition ” Prog. Theor. Phys., Vol. 28 No. 4 (Oct. 1962) 757-759.

55. M. Taketani, Y. Fujimoto and K. Yokoi Proceedings of the Fifth Interamerican Seminar on Cosmic Rays, La Paz, Bolivia, Vol. II (1962), LIV.

56. M. Taketani and T. Tati Neutrino Unified Model Based on Two Kinds of Neutrinos Prog. Theor. Phys., Vol. 30 No. 1 (Jul. 1963) 134-136.

57. K. Aizu, Y. Fujimoto, H. Hasegawa, S. Hayakawa and M. Taketani A Model of Star-like Objects Prog. Theor. Phys., Vol. 32 No. 6 (Dec. 1964) 971-973.

58. K. Aizu, Y. Fujimoto, H. Hasegawa, K. Kawabata and M. Taketani A New Diagram Representing the Evolution of Galaxies Prog. Theor. Phys., Vol. 32 No. 6 (Dec. 1964) 973-975.

59. Y. Fujimoto, H. Hasegawa and M. Taketani Extra-Galactic Cosmic Rays Prog. Theor. Phys., Vol. 32 No. 6 (Dec. 1964) 975-977.

60. M. Taketani Preface Origin of Cosmic Rays Prog. Theor. Phys. Suppl., No. 30 (1964) 1 (pub. date: Apr. 1965).

61. K. Aizu, Y. Fujimoto, H. Hasegawa and M. Taketani Introduction and General Review Origin of Cosmic Rays Prog. Theor. Phys. Suppl., No. 30 (1964) 2-31 (pub. date: Apr. 1965).

62. Y. Fujimoto, H. Hasegawa and M. Taketani Part 1. Cosmic Rays in Galactic and Extragalactic Space Origin of Cosmic Rays Prog. Theor. Phys. Suppl., No. 30 (1964) 32-85 (pub. date: Apr. 1965).

63. M. Taketani, Y. Fujimoto and K. Yokoi Appendices. Evolution of Galaxies and Its Relation to Origin of Cosmic Rays (reprinted from Proc. Fifth Interamerican Seminars on Cosmic Rays, La Paz (Bolivia), 1962, Vol. II) Origin of Cosmic Rays Prog. Theor. Phys. Suppl., No. 30 (1964) 244-249 (pub. date: Apr. 1965).

64. M. Taketani Preface Structure and Evolution of Galaxies Prog. Theor. Phys. Suppl., No. 31 (1964) 1 (pub. date: Mar. 1965).

65. T. Hatanaka, S. Hayakawa, K. Ishida and M. Taketani Part I. Evolution of Galaxies - Development of Its Research in Japan - Structure and Evolution of Galaxies Prog. Theor. Phys. Suppl., No. 31 (1964) 2-34 (pub. date: Mar. 1965).

66. K. Aizu, Y. Fujimoto, H. Hasegawa, K. Kawabata and M. Taketani Part II. Radio Galaxies Structure and Evolution of Galaxies Prog. Theor. Phys. Suppl., No. 31 (1964) 35-76 (pub. date: Mar. 1965).

67. M. Taketani, T. Hatanaka and S. Obi Part VI. Appendix. Populations and Evolution of Stars Structure and Evolution of Galaxies (reprinted from Prog. Theor. Phys., Vol. 15 (1956) 89-94) Structure and Evolution of Galaxies Prog. Theor. Phys. Suppl., No. 31 (1964) 156-161 (pub. date: Mar. 1965).

68. K. Aizu, Y. Fujimoto, H. Hasegawa and M. Taketani An Evolutional Track of Galaxies in L − (L/V) Diagram Prog. Theor. Phys., Vol. 33 No. 5 (May 1965) 996-997.

69. Y. Katayama and M. Taketani Neutrino Unified Model for Elementary Particles Prog. Theor. Phys., Vol. 33 No. 6 (Jun. 1965) 1141-1143.

70. M. Taketani and Y. Fujimoto High Energy Nuclear Interactions and Structure of Elementary Particles Commemoration Issue for the Thirtieth Anniversary of the Meson Theory by Dr. H. Yukawa Prog. Theor. Phys. Suppl., Extra No. (Sep. 1965) 651-660.

71. M. Taketani On the Meson Theory of Nuclear Forces Proceedings of the International Conference on Elementary Particles, Kyoto (1965) 170-181 (pub. date: 1966).

72. M. Taketani, S. Nakamura and Y. Fujimoto On the Fundamental Concepts in the Theory of Elementary Particles. I

- History of the Tomonaga Theory of Fields - Dedicated to Professor Sin-itiro Tomonaga on the Occasion of his Sixtieth Birthday Prog. Theor. Phys. Suppl., No. 37 & 38 (Dec. 1966) 433-439.

73. M. Taketani, S. Nakamura, Y. Fujimoto, M. Nagasaki, S. Furuichi and S. Sato On the Fundamental Concepts in the Theory of Elementary Particles. II

- Present Problems - Dedicated to Professor Sin-itiro Tomonaga on the Occasion of his Sixtieth Birthday Prog. Theor. Phys. Suppl., No. 37 & 38 (Dec. 1966) 440-450.

74. M. Nagasaki, T. Hirasawa and M. Taketani On the Analysis of Nucleon-Nucleon Scattering with Large Momentum Transfers Prog. Theor. Phys. Vol. 38 No. 2 (Aug. 1967) 514-516.

75. M. Taketani Introduction - Our History and the Underlying Philosophy - Part I. Nuclear Force in Dynamical Region Nuclear Forces Prog. Theor. Phys. Suppl., No. 39 (Aug. 1967) 1-22.

76. T. Hirasawa, K. Aizu and M. Taketani Formation of Galaxies from Hydrogen Gas Prog. Theor. Phys., Vol. 41 No. 3 (Mar. 1969) 835-838.

77. K. Aizu and M. Taketani Radio Evolution of Pulsars Prog. Theor. Phys., Vol. 44 No. 2 (Aug. 1970) 363-374.

78. M. Nagasaki and M. Taketani Level Structure of Hadrons Prog. Theor. Phys., Vol. 44 No. 4 (Oct. 1970) 1112-1114.

79. K. Aizu, S. Hayakawa, H. Sato and M. Taketani Chapter I. General Review and Summary - Astrophysical Cosmology - Evolution of the Universe and Formulation of Galaxies Prog. Theor. Phys. Suppl., No. 49 (1971) 1-10 (pub. date: Feb. 1972).

80. K. Aizu, H. Tabara and M. Taketani Chapter VII. Quasistellar Objects Evolution of the Universe and Formulation of Galaxies Prog. Theor. Phys. Suppl., No. 49 (1971) 228-247 (pub. date: Feb. 1972).

81. M. Taketani Methodological Approaches in the Development of the Meson Theory of Yukawa in Japan Historical Introduction Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 12-24 (pub. date: Mar. 1972).

82. M. Taketani Dialectics of Nature - On Quantum Mechanics - Part I Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 27-36 (pub. date: Mar. 1972).

83. M. Taketani A New Stage in the Development of Modern Physics

- On Research Contribution of Nobelians, the Joliot-Curie’s - Part I Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 37-48 (pub. date: Mar. 1972).

84. M. Taketani On Research Contribution of Niels Bohr Part I Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 49-52 (pub. date: Mar. 1972).

85. M. Taketani On Formation of the Newton Mechanics Part I Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 53-64 (pub. date: Mar. 1972).

86. M. Taketani Observation Problem of Quantum Mechanics Part I Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 65-72 (pub. date: Mar. 1972).

87. M. Taketani Appendix. (1) Light Quanta and Special Relativity Part I Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 73-76 (pub. date: Mar. 1972).

88. M. Taketani Appendix. (2) Substance and Essence Part I Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., 50 (1971) 77-80 (pub. date: Mar. 1972).

89. M. Taketani Conflict between Matter and Field

- An Analysis of the Difficulties of the Theory of Elementary Particles - Part II Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 89-97 (pub. date: Mar. 1972).

90. M. Taketani New Development of Theory of Elementary Particles

- Theory of Shoichi Sakata and Related Problems - Part II Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 127-139 (pub. date: Mar. 1972).

91. M. Taketani, S. Nakamura, K. Ono and M. Sasaki Appendix. Two Meson Theory Part II Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 140-143 (pub. date: Mar. 1972).

92. M. Taketani and S. Nakamura On the Method of the Theory of Elementary Particles Part III Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl. No. 50 (1971) 145-149 (pub. date: Mar. 1972).

93. M. Taketani Program of Research on Nuclear Force (1953 Proposal) Part III Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl. No. 50 (1971) 159-162 (pub. date: Mar. 1972).

94. M. Taketani Appendix. (3) A Thought on Structure of Elementary Particles Part IV Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 214-219 (pub. date: Mar. 1972).

95. M. Taketani Stellar Evolution and Galactic Evolution

- From THO Theory to Star-like Problem - Part V. Towards New Concepts of Nature Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 220-230 (pub. date: Mar. 1972).

96. M. Taketani and S. Nakamura Nuclear Physics and Astronomy Part V. Towards New Concepts of Nature Philosophical and Methodological Problems in Physics Prog. Theor. Phys. Suppl., No. 50 (1971) 231-234 (pub. date: Mar. 1972).

97. M. Nagasaki and M. Taketani Excitation of Hadronic Matter and H -Quantum Prog. Theor. Phys., Vol. 49 No. 1 (Jan. 1973) 368-370.

98. K. Aizu and M. Taketani On the Phase of Pulsar Evolution

- A Comment on Tsarevsky’s Paper - Prog. Theor. Phys., Vol. 49 No. 4 (Apr. 1973) 1371-1372.

99. M. Nagasaki and M. Taketani Angular Distribution of p-p Scattering at High Energies and Radius of the Hard Core Prog. Theor. Phys., Vol. 49 No. 5 (May 1973) 1768-1770.

100. M. Nagasaki and M. Taketani Angular Distribution of p-p Elastic Scattering at High Energies and Distribution of Absorption Prog. Theor. Phys., Vol. 49 No. 6 (Jun. 1973) 2145-2147.

101. M. Taketani, T. Tati, K. Mori, M. Nagasaki, Y. Fujimoto and S. Hasegawa Introduction and Summary

- Theoretical Models for Multiple Production of Mesons - Ultra-High Energy Phenomena and New State of Matter Prog. Theor. Phys. Suppl., No. 54 (1973) 1-16 (pub. date: Jun. 1974).

102. M. Nagasaki and M. Taketani Angular Distribution of p-p Elastic Scattering at High Energy and Radius of the Hard Core Ultra-High Energy Phenomena and New State of Matter Prog. Theor. Phys. Suppl., No. 54 (1973) 51-72 (pub. date: Jun. 1974).

103. M. Nagasaki and M. Taketani Excitation of Hadronic Matter and H -quantum Ultra-High Energy Phenomena and New State of Matter Prog. Theor. Phys. Suppl., No. 54 (1973) 73-92 (pub. date: Jun. 1974).

104. M. Taketani Opening Door to New Physics

- History and Underlying Philosophy of the Research Group on Ultra-High Energy Phenomena and New State of Matter - An attached paper with 14 pages to Prog. Theor. Phys. Suppl., No. 54 (1973) (pub. date: Jun. 1974).

105. M. Taketani and M. Nagasaki “ Centauro ” Phenomenon and the Decay of UH -Quantum Prog. Theor. Phys., Vol. 53 No. 4 (Apr. 1975) 1212-1213.

106. M. Taketani On Present-Day Problems in Theory of Elementary Particles Fire-Balls in Pion Multiple Production Prog. Theor. Phys. Suppl., No. 76 (Nov. 1983) 224-242.

107. Mituo Taketani Theoretical Study of Cosmic Rays in Japan Early History of Cosmic Ray Studies: Personal Reminiscences with Old Photographs (Astrophysics and Space Science Library ; v. 118) ed. by Yataro Sekido and Harry Elliot, D. Reidel, (1985) 285-294.

108. H. Yukawa, S. Sakata and M. Taketani Appendix. On the Theory of the New Particle in Cosmic Ray Prog. Theor. Phys. Suppl., No. 85 (1985) 310-311 (pub. date: Mar. 1986) (Published as the Proceedings of the Kyoto International Symposium, The Jubilee of the Meson Theory, Kyoto, Japan, August 15-17,1985) An unpublishecd paper sent to the Physical Review on 4th October 1937.

109. M. Taketani, L. M. Brown, Y. Fujimoto, M. Konuma, M. Nagasaki and T. Tsuji 7. Physics and Philosophy - Interview with Mituo Taketani - II. Interviews Part I. Proceedigns of the Japan-USA Collaborative Workshops on the History of Particle Theory in Japan, 1930-1950 (Japan-USA Collaboration September 1978-July 1979) Elementary Particle Theory in Japan, 1930-1960 Proceedigns of the Japan-USA Collaborative Workshops Prog. Theor. Phys. Suppl., No. 105 (Dec. 1991) 86-98.

参考文献

[1] 矢野 忠, 武谷三男博士の業績リスト, 素粒子論研究, 第 107 巻, 第 4 号 (2003.7) 9-18

[2] 矢野 忠, 武谷三男博士の業績リスト (第 2 版), 素粒子論研究, 第 116 巻, 第 5 号 (2008.12) 159-168

[3] 矢野 忠, 武谷三男博士の著作目録, 素粒子論研究, 第 109 巻, 第 2 号 (2004.5) 16-33

[4] 矢野 忠, 武谷三男博士の業績・著作補遺, 素粒子論研究, 第 111 巻, 第 2 号 (2005.5) 100-103

[5] 矢野 忠, 武谷三男博士の業績・著作補遺 2, 素粒子論研究, 第 113 巻, 第 1 号 (2006.4) 1-2

[6] 矢野 忠, 武谷三男博士の著作目録 (第 2 版), 素粒子論研究, 第 116 巻, 第 5 号 (2008.12) 169-187

[7] 矢野 忠, 武谷三男博士の著作目録 (第 3 版), 素粒子論研究電子版，投稿中